# A Unified Dynamical Field Theory of Learning, Inference, and Emergence

### v1

@@ B. Learning as Reorganization of the Time-Scale Density of States
A particularly important outcome of this work is the identification of the time-scale density of states (TDOS) as a physically meaningful descriptor of cognition. Neural systems are composed of heterogeneous, irregularly connected, and intrinsically stochastic elements, a fact that has long obscured how stable inference and memory can arise. Within the present framework, this apparent paradox is resolved by recognizing that cognitive function is governed not by microscopic precision, but by the collective organization of dynamical time scales. Learning and homeostatic adaptation reshape the TDOS, selectively stabilizing slow collective modes while suppressing unstable or spurious directions. Cognition thus corresponds to the controlled persistence of these slow modes, which integrate information over extended time intervals and remain robust against noise and microscopic variability.
@@
From the perspective of artificial intelligence, the present theory suggests that the success of modern architectures does not primarily stem from architectural complexity, but from their implicit approximation of the unified dynamical equation. Residual connections, normalization layers, and attention mechanisms can be understood as discrete implementations of continuous recurrent flows constrained by learned geometry. This viewpoint points toward principled strategies for architecture design, emphasizing stability, reentry, and geometric regularization over ad hoc complexity.
@@
We have developed a unified dynamical field-theoretic framework for learning, inference, and emergence, identifying a minimal physical principle underlying cognitive dynamics. Within this framework, inference corresponds to saddle-point trajectories of an MSRJD action, while fluctuation-induced loop corrections generically generate collective modes and characteristic time scales beyond mean-field descriptions.
A central result of this work is that cognitive function is governed not by microscopic units or precise activity patterns, but by the collective organization of dynamical time scales. We introduced the time-scale density of states as a compact and physically meaningful descriptor of this organization, providing a unified diagnostic for inference dynamics across biological and artificial systems. Learning and homeostatic regulation are naturally interpreted as processes that reshape the TDOS, selectively stabilizing slow collective modes that support robust inference, memory, and context-dependent computation.
By shifting the focus from individual components to collective time-scale structure, the present theory explains how stable cognition can emerge from heterogeneous, stochastic, and irregular substrates. More broadly, it suggests that learning, inference, and emergence are not distinct processes, but manifestations of a single dynamical principle governing high-dimensional adaptive systems.
@@ Appendix B Appendix B: Derivation of the FHRN Decomposition
Projecting $\bm{y}$ onto the dominant eigenspace of $\bm{W}$, $\bm{y}\approx r\,\bm{v}_{0}$, yields
$\bm{y}^{\top}\bm{W}\bm{y}\approx\rho\,r^{2},$ (S18)
and hence the exact radial equation
@@ Appendix C Appendix C: Emergent Riemannian Geometry from Homeostatic Shell Dynamics
Homeostasis strongly suppresses radial excursions while leaving tangential degrees of freedom comparatively unconstrained. The resulting effective continuous-time dynamics is therefore most naturally expressed in terms of an anisotropic, state-dependent metric of the form (S26). Importantly, this metric is not imposed externally. Rather, it provides a compact representation of the coexistence of strong radial stabilization and tangential freedom induced by homeostatic shell dynamics.
(I) Non-Euclidean nature of the induced geometry: A crucial observation is that the metric $G(y)$ is not spatially constant and therefore does not correspond to a Euclidean geometry. Since both $P_{r}(y)$ and $P_{\perp}(y)$ depend explicitly on the state vector $y$, their derivatives are generically nonzero, implying
@@
throughout the shell neighborhood.
@@
does not vanish even arbitrarily close to the stable shell. This immediately implies a nontrivial Riemann curvature tensor,
@@
Thus, the shell formed by homeostatic regulation is not merely a constrained subset of an ambient Euclidean space. Instead, it induces a genuinely non-Euclidean state-space geometry endowed with intrinsic Riemann curvature.
(II) Geodesic deviation and cognitive tidal forces: The dynamical significance of this curvature becomes apparent through the geodesic deviation equation. Let $\zeta^{i}(t)$ denote the separation vector between two nearby inference trajectories. Its evolution obeys
@@
If the sectional curvature associated with the relevant two-dimensional plane is positive, nearby trajectories undergo focusing and converge. If the sectional curvature is negative, trajectories defocus and branch. Crucially, this behavior arises independently of the Hessian of the potential $\Phi$, external noise, or instability of initial conditions. It represents a purely geometric effect originating from the curvature of the state-space metric.
@@
The continuous-time limit of the FHRN dynamics exhibits the following structure. Homeostatic regulation stabilizes activity on a finite-radius shell. This stabilization enforces a radial–tangential anisotropy in the effective dynamics, which is naturally encoded by a state-dependent metric. As a direct consequence, a non-Euclidean Riemannian geometry emerges inevitably in the state space.
@@
This geometric structure regulates inference sensitivity, trajectory branching, and the selection of slow collective modes in ways that cannot be captured by potential structure alone. We interpret these curvature-induced effects as cognitive tidal forces, providing a purely geometric mechanism for sensitivity and diversification in inference dynamics.

### v2

@@ B. Learning as Reorganization of the Time-Scale Density of States
Cognitive field interpretation.
The geometric structure underlying the unified dynamical equation naturally admits an interpretation in terms of a cognitive field. By this term, we do not posit a new physical field in the electromagnetic or relativistic sense, but refer to an effective dynamical field defined on the collective state space, whose structure jointly governs learning, inference, and emergent cognitive dynamics.
Within this interpretation, the collective state evolves under a state-space vector field whose gradient component, shaped by the effective potential $\Phi$ and the metric $G$, implements inference as relaxation toward dynamically stabilized manifolds, while the non-conservative reentrant component defines an autonomous solenoidal flow that redistributes activity along these manifolds without altering the potential. Crucially, the presence of this reentrant component is not a modeling embellishment but a geometric necessity once a metric structure on state space is specified.
Interpreting the unified dynamics as a cognitive field places learning, inference, and emergence on the same conceptual footing as field-theoretic descriptions in statistical physics, where macroscopic behavior is governed by effective fields rather than microscopic details. In this sense, the cognitive field provides a compact and principled language for summarizing how learned geometry, reentrant dynamics, and collective time-scale organization jointly support stable yet flexible cognition in high-dimensional, noisy systems.
VII.2 B. Collective Time Scales and the Time-Scale Density of States
A particularly important outcome of this work is the identification of the time-scale density of states as a physically meaningful descriptor of cognition. Neural systems are composed of heterogeneous, irregularly connected, and intrinsically stochastic elements, a fact that has long obscured how stable inference and memory can arise. Within the present framework, this apparent paradox is resolved by recognizing that cognitive function is governed not by microscopic precision, but by the collective organization of dynamical time scales. Learning and homeostatic adaptation reshape the TDOS, selectively stabilizing slow collective modes while suppressing unstable or spurious directions through the joint modulation of potential structure and state-space geometry. Cognition thus corresponds to the controlled persistence of these slow modes, which integrate information over extended time intervals and remain robust against noise and microscopic variability.
@@
VII.3 C. Implications for Artificial Intelligence and Architecture Design
In the context of artificial intelligence, the present theory suggests that the empirical success of modern architectures does not primarily stem from architectural complexity, but from their implicit approximation of a unified underlying dynamical equation. Residual connections, normalization layers, and attention mechanisms may be interpreted as discrete implementations of continuous recurrent flows constrained by a learned state-space geometry. From this standpoint, principled strategies for architecture design emerge naturally, emphasizing stability, reentry, and geometric regularization rather than ad hoc structural elaboration.
Within this dynamical interpretation, the expressive power of modern transformer architectures admits a clarifying explanation. Their apparent intelligence does not arise solely from the explicit optimization of an objective function, but from the ability to dynamically reshape the effective geometry of representation space in a state-dependent manner. Attention mechanisms implement this adaptive geometric reweighting at the level of computation, while the training objective acts as a background potential that biases selection within the induced geometry. Accordingly, attention does not introduce a fundamentally new mode of computation, but realizes a discrete approximation to inference on a learned, context-dependent state-space geometry.
VII.4 D. Outlook and Open Questions
@@
We have developed a unified dynamical field-theoretic framework for learning, inference, and emergence, identifying a minimal physical principle underlying cognitive dynamics. Within this framework, inference corresponds to saddle-point trajectories of an MSRJD action, realized as deterministic relaxation flows on a learned, state-dependent geometry of collective state space, while fluctuation-induced loop corrections generically generate collective modes and characteristic time scales beyond mean-field descriptions.
A central result of this work is that cognitive function is governed not by microscopic units or precise activity patterns, but by the collective organization of dynamical time scales across the full joint state space. We introduced the time-scale density of states as a compact and physically meaningful descriptor of this organization, providing a unified diagnostic for inference dynamics across biological and artificial systems. Learning and homeostatic regulation are naturally interpreted as processes that reshape the effective potential and the underlying state-space geometry, thereby reorganizing the TDOS and selectively stabilizing slow collective modes that support robust inference, memory, and context-dependent computation.
By shifting the focus from individual components to collective time-scale structure, and to the geometry that governs sensitivity and stability of inference trajectories, the present theory explains how stable cognition can emerge from heterogeneous, stochastic, and irregular substrates. More broadly, it suggests that learning, inference, and emergence are not distinct processes, but manifestations of a single dynamical principle governing high-dimensional adaptive systems.
@@ Appendix B Appendix B: Derivation of the FHRN Decomposition
Projecting $\bm{y}$ onto the dominant eigenspace of $\bm{W}$ and assuming alignment with the leading eigenvector, $\bm{y}\approx r\,\bm{v}_{0}$, one obtains
$\bm{y}^{\top}\bm{W}\bm{y}\approx\rho\,r^{2}.$ (S18)
Under this dominant–eigenmode approximation, the radial dynamics closes as
@@ Appendix C Appendix C: Emergent Riemannian Geometry from Homeostatic Shell Dynamics
Homeostatic regulation strongly suppresses radial excursions of collective activity while leaving tangential degrees of freedom comparatively unconstrained. As a result, the effective continuous-time dynamics near the stabilized shell is most naturally expressed in terms of an anisotropic, state-dependent metric of the form Eq. (S26). Importantly, this metric is not imposed externally. Rather, it provides a compact representation of the differential sensitivity of the underlying probabilistic inference dynamics, and may be interpreted as an effective Fisher information metric induced by homeostatic shell dynamics and pulled back to the collective state space.
(I) Non-Euclidean nature of the induced geometry: A crucial consequence of this construction is that the metric $G(y)$ is not spatially constant and therefore does not correspond to a Euclidean geometry. Since both the radial and tangential projectors, $P_{r}(y)$ and $P_{\perp}(y)$, depend explicitly on the state vector $y$, their derivatives are generically nonzero, implying
@@
throughout the neighborhood of the homeostatic shell.
@@
does not vanish even arbitrarily close to the stabilized manifold. This immediately implies a nontrivial Riemann curvature tensor,
@@
Thus, the shell formed by homeostatic regulation is not merely a constrained subset of an ambient Euclidean space. Instead, it induces a genuinely non-Euclidean state-space geometry endowed with intrinsic Riemann curvature, reflecting the state-dependent sensitivity of probabilistic inference.
(II) Geodesic deviation and cognitive tidal forces: The dynamical implications of this curvature become explicit through the geodesic deviation equation. Let $\zeta^{i}(t)$ denote the separation vector between two nearby inference trajectories evolving under identical dynamics. Its evolution obeys
@@
Depending on the local sectional curvature of the state space, nearby trajectories may undergo focusing or defocusing. Positive curvature leads to convergence of trajectories, while negative curvature induces divergence and branching. Crucially, these effects arise independently of the Hessian of the potential $\Phi$, stochastic fluctuations, or instabilities of initial conditions. They represent a purely geometric mechanism originating from the curvature of the state-space metric.
@@
In the continuous-time limit of the FHRN dynamics, homeostatic regulation stabilizes activity on a finite-radius shell, enforcing a pronounced radial–tangential anisotropy in the effective dynamics. This anisotropy is naturally encoded in the induced state-dependent metric and inevitably gives rise to a non-Euclidean Riemannian geometry.
@@
We interpret the resulting curvature-induced relative acceleration between nearby inference trajectories as cognitive tidal forces: a geometric manifestation of the information-theoretic sensitivity encoded in the Fisher metric, regulating inference sensitivity, trajectory branching, and the selection of slow collective modes beyond what can be captured by potential structure alone.
Appendix D Appendix D: A minimal Fisher-information derivation of the shell metric
Here we show, in a concrete and minimal probabilistic setting, how a Fisher-information metric on the collective state space naturally acquires the same radial–tangential projector structure as the FHRN metric, $G(y)=g_{r}(r)P_{r}+g_{\perp}(r)P_{\perp}$. The purpose is not to claim a unique generative model, but to exhibit a simple class of distributions whose local information geometry necessarily takes this form whenever radial and tangential sensitivities are separated.
(I) Anisotropic Gaussian family on $\mathbb{R}^{N}$: Let $x\in\mathbb{R}^{N}$ denote the latent collective state and $z\in\mathbb{R}^{N}$ an “internal observation” (or self-generated prediction target) drawn from a Gaussian likelihood centered at $x$,
$p(z|x)=\mathcal{N}\!\big(z;\,x,\Sigma(x)\big),$ (S34)
with an anisotropic covariance aligned with radial and tangential subspaces at the current state:
$\Sigma(x)=\sigma_{r}^{2}(r)\,P_{r}(x)+\sigma_{\perp}^{2}(r)\,P_{\perp}(x),\qquad r=\|x\|,$ (S35)
$\qquad P_{r}=\frac{xx^{\top}}{r^{2}},\quad P_{\perp}=I-P_{r}.$
This is the simplest rotationally covariant way to encode the idea that radial fluctuations are more strongly suppressed than tangential ones near a stabilized shell (e.g., $\sigma_{r}(r)\ll\sigma_{\perp}(r)$ for $r\approx r_{\ast}$).
Since $P_{r}$ and $P_{\perp}$ are orthogonal projectors, the inverse covariance has the same projector form,
$\Sigma^{-1}(x)=\frac{1}{\sigma_{r}^{2}(r)}\,P_{r}(x)+\frac{1}{\sigma_{\perp}^{2}(r)}\,P_{\perp}(x).$ (S36)
(II) Fisher information for the mean parameter: For a Gaussian with mean $x$ and (locally) fixed covariance, the score function is
$\partial_{i}\log p(z|x)=\big[\Sigma^{-1}(x)\,(z-x)\big]_{i}\quad\text{(neglecting $\partial_{i}\Sigma$ terms)}.$ (S37)
This approximation is appropriate when the dominant sensitivity arises from mean shifts (the “prediction error” zxz-x) and the variation of $\Sigma$ with $x$ is subleading on the scale of typical fluctuations. Under this standard simplification,
$F_{ij}(x)\equiv\mathbb{E}_{z|x}\!\Big[\partial_{i}\log p(z|x)\;\partial_{j}\log p(z|x)\Big]=\big[\Sigma^{-1}(x)\big]_{ij},$ (S38)
because $\mathbb{E}_{z|x}\big[(z-x)(z-x)^{\top}\big]=\Sigma(x)$.
Combining Eqs. (S36)–(S38) yields the Fisher metric in projector form:
$F(x)=f_{r}(r)\,P_{r}(x)+f_{\perp}(r)\,P_{\perp}(x),\qquad f_{r}(r)=\sigma_{r}^{-2}(r),\quad$ (S39)
$f_{\perp}(r)=\sigma_{\perp}^{-2}(r).$
Thus, whenever radial and tangential uncertainties differ, the induced information geometry is spherically symmetric but anisotropic: it preserves rotational covariance while assigning distinct local length scales to radial and tangential directions.
(III) Identification with the FHRN shell metric: Equation (S39) matches the metric structure used in the FHRN decomposition upon the identification
$G(x)\;\propto\;F(x),\qquad g_{r}(r)\propto f_{r}(r)=\sigma_{r}^{-2}(r),\qquad$ (S40)
$g_{\perp}(r)\propto f_{\perp}(r)=\sigma_{\perp}^{-2}(r).$
In particular, homeostatic stabilization of the shell corresponds to $\sigma_{r}(r_{\ast})\ll\sigma_{\perp}(r_{\ast})$ and therefore to a strongly anisotropic Fisher geometry with $g_{r}(r_{\ast})\gg g_{\perp}(r_{\ast})$, capturing the empirical fact that radial excursions are tightly constrained while tangential degrees of freedom remain comparatively free.
If one retains the full $x$-dependence of $\Sigma(x)$, additional terms from $\partial_{i}\Sigma$ enter the score and Fisher information. However, for rotationally covariant families of the form Eq. (S35), these corrections preserve the same radial–tangential decomposition and merely renormalize the scalar coefficients $f_{r}(r),f_{\perp}(r)$. Therefore the projector structure is robust: it follows from symmetry and the separation of radial and tangential sensitivities, not from the Gaussian choice itself.